Circular No. 6449

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 P2 (RUSSELL-WATSON)

The following CCD precise positions have been reported:

1996 UT α_{2000} δ_{2000} m_{1} Observer
Aug. 12.57184 1^{h}40^{m}50.^{s}85 −30^{o}10′29′′.9  Garradd
12.57484 1 40 50.92 −30 10 31.4  _{”}
12.70747 1 40 52.44 −30 11 57.7 13.7 Abe
12.71858 1 40 52.57 −30 12 04.7  _{”}
12.72473 1 40 52.63 −30 12 08.5 14.1 Ikari
12.73673 1 40 52.75 −30 12 15.8 14.2  _{”}
13.60493 1 41 02.95 −30 21 00.1  Garradd
13.60627 1 41 02.94 −30 21 00.9  _{”}

G. J. Garradd (Loomberah). 0.25-m f/4.1 reflector. First image involved

with flaw. H. Abe (Yatsuka). 0.26-m f/6.0 reflector. Coma diameter 20′′, tail ∼ 5′

long. Communicated by S. Nakano. Y. Ikari (Moriyama). 0.25-m f/6.3 reflector. Strong condensation with tail

in p.a. 250^{o}. Communicated by S. Nakano.

The parabolic orbital elements below are from 11 observations Aug. 10–13. It is possible that the comet is of short period, several contributors remarking on a general similarity of the orbit to that of comet P/1994 J3.

T = 1996 Feb. 21.064 TT  ω = 187.^{o}526
Ω = 103.431 }2000.0
q = 2.07725 AU  i = 29.725
1996 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 5 1^{h}38.^{m}47 −28^{o}53.^{′}9 2.234 2.829 115.^{o}9 18.^{o}8 12.8
10 1 40.21 −29 44.4 2.232 2.866 119.2 18.0 12.8
15 1 41.27 −30 35.2 2.234 2.904 122.3 17.1 12.9
20 1 41.64 −31 25.4 2.239 2.943 125.4 16.3 12.9
25 1 41.34 −32 14.2 2.247 2.981 128.3 15.4 13.0
30 1 40.39 −33 00.7 2.260 3.020 131.0 14.6 13.1
Sept. 4 1 38.84 −33 44.0 2.278 3.059 133.4 13.9 13.1
9 1 36.74 −34 23.2 2.300 3.099 135.5 13.2 13.2
14 1 34.14 −34 57.6 2.326 3.139 137.1 12.6 13.3

Total visual magnitude estimate by A. Hale, Cloudcroft, NM (0.41-m reflector): Aug. 13.43 UT, 12.7 (coma diameter 1′.3).

1996 August 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden